OpenStudy (anonymous):

please help me with this trig.. sinu=3/5 and sinv=4/5, evaluate sin(u-v) and tan(u+v)

OpenStudy (anonymous):

Firstly you should calculate cos(u) and cos(v). Do you know how you will calculate these?

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

See, we have an identity which says: \[\sin^2(x) + \cos^2(x) = 1\] Here let us find out the value of cos(x). \[\cos^2(x) = 1- \sin^2(x)\] \[\implies \cos(x) = \sqrt{1 - \sin^2(x)}\] Getting this much?

OpenStudy (anonymous):

oh ok i can do that, then do i use sin(u-v)=sinucosv-cosusinv?

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

Firstly find cos(u) and cos(v) by using the formula I wrote above.. Then use that formula you wrote..

OpenStudy (anonymous):

thank you very much friend, you're the best

OpenStudy (anonymous):

You are welcome dear.. And I can be better but I am not the best..